if the temperature of the helium balloon were increased from 30 celcius to 35 celcius and the volume of the ballon only expanded from 0.47 L to 1.55 L, how much pressure would be inside the ballon?
1
Expert's answer
2020-04-22T09:18:58-0400

According to the combined gas law equation:

p1V1/T1 = p2V2/T2

where p1 - initial pressure, V1 - initial volume, T1 - initial temperature, p2 - final pressure, V2 - final volume, T2 - final temperature.

From here:

p2/p1 = V1T2/V2T1

As 30°C = 303 K and 35°C = 308 K:

p2/p1 = 0.47 L × 308 K / 1.55 L × 303 K = 0.31

As the value is below zero, the final pressure is lower than the initial pressure. After conversion:

p1/p2 = 1 / 0.31 = 3.2

Answer: The pressure in the ballon is 3.2 times lower than to the initial pressure.